### v1

@@ 3. Regularity of self-similar measures
We now bound $I_{3}$. Using that $\sigma$ has a finite moment (Lemma˜3.1), and the large deviation principle for $(r_{\phi})_{\phi\sim\lambda^{*n}}$, we find for some $c_{3}=c_{3}(\sigma)>0$,
$\lambda^{*n}\{\phi\,:\,\phi_{\star}\sigma(\mathbb{R}^{d}\smallsetminus B_{\varepsilon^{1/2}}(\phi(0)))\geq\varepsilon^{c_{3}}\}\ll\varepsilon^{c_{3}},$
whence
$I_{3}\leq$ $\int_{\phi(0)\in B_{R}\smallsetminus F_{\delta}}\phi_{\star}\sigma[E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))]\,\mathrm{d}\lambda^{*n}(\phi)+O(\varepsilon^{c_{3}}).$
Consider $\phi$ as in the above integral, and $\boldsymbol{s}\in E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))$. By Taylor expansion, and the fact that $\|\cdot\|_{C^{2}(B_{R+1})}\ll_{l}R^{l}\|\cdot\|$ on ${\mathcal{P}}_{d,l}$, the conditions $\|P\|=1$ and $\|\boldsymbol{s}-\phi(0)\|<\varepsilon^{1/2}$ imply
$|P(\boldsymbol{s})-P(\phi(0))-\big{\langle}\nabla P(\phi(0)),\,\,\boldsymbol{s}-\phi(0)\big{\rangle}|\ll_{l}\varepsilon R^{l}.$
As $|P(\boldsymbol{s})|\leq\varepsilon$, setting $v_{P,\phi}:=P(\phi(0))-\big{\langle}\nabla P(\phi(0)),\,\,\phi(0)\big{\rangle}$, we deduce
$|\big{\langle}\nabla P(\phi(0)),\,\,\boldsymbol{s}\big{\rangle}-v_{P,\phi}|\ll_{l}\varepsilon R^{l}$
But $\|\nabla P(\phi(0))\|>\delta$ by assumption, whence $\boldsymbol{s}$ belongs to the $O_{l}(\varepsilon R^{l}\delta^{-1})$-neighborhood of an affine hyperplane. Applying $\phi^{-1}$, and the non-concentration near affine subspaces from Proposition˜3.2, we obtain
$\phi_{\star}\sigma[E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))]\ll_{l}(\varepsilon R^{l}\delta^{-1}r_{\phi}^{-1})^{c_{4}}$
where $c_{4}=c_{4}(\sigma)>0$. Using the large deviation principle for $(r_{\phi})_{\phi\sim\lambda^{*n}}$, we may integrate to obtain
(3.5) $I_{3}\ll_{l}(\varepsilon^{1/8}R^{l}\delta^{-1})^{c_{4}}+\varepsilon^{c_{3}}.$
In the end, combining (3.2), (3.4), (3.5), and choosing $\delta=\varepsilon^{1/16}$, $R=\varepsilon^{-\alpha}$ with $\alpha\lll_{\sigma,l}1$, we have proven that for $c^{\prime}\lll_{\gamma,c_{1},\dots,c_{4}}1$,
$\sigma(E_{\varepsilon})\ll\varepsilon^{c^{\prime}}.$
@@
Let $k,l\geq 1$. Let $P:\mathbb{R}^{dk}\rightarrow\mathbb{R}$ be a polynomial map of degree at most $l$ and such that $\|P\|\geq 1$. Then for every $\varepsilon>0$,
$\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\}\leq C\varepsilon^{c}$
where $C=C(\sigma,k,l)>1$ and $c=c(\sigma,l)>0$.
@@
We argue by induction on $k$. The case $k=1$ is ˜3.3. Now given $k\geq 2$, we assume the result holds up to k1k-1 and prove it for $k$. Consider $Q:\mathbb{R}^{d}\rightarrow{\mathcal{P}}_{d(k-1),l},\boldsymbol{s}\mapsto P(\boldsymbol{s},\,\cdot\,)$, which is a polynomial map whose coordinates have degree at most $l$. Moreover, endowing ${\mathcal{P}}_{d(k-1),l}$ with the standard basis, $Q$ and $P$ have same coefficients, in particular $\|Q\|=\|P\|\geq 1$. It follows from the $k=1$ case that for some $c_{1}=c_{1}(\sigma,l)>0$, for every $\delta>0$,
$\sigma\{\boldsymbol{s}\in\mathbb{R}^{d}\,:\,\|Q(\boldsymbol{s})\|\leq\delta\}\ll_{l}\delta^{c_{1}}.$
We deduce
$\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\}$
$\leq$ $\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\text{ and }\|Q(\boldsymbol{s}_{1})\|>\varepsilon^{1/2}\}+O(\varepsilon^{c_{1}/2})$
$\ll$ $_{k,l}\varepsilon^{c_{2}/2}+\varepsilon^{c_{1}/2},$
where the last inequality follows from the induction hypothesis with parameter k1k-1, and exponent $c_{2}=c_{2}(\sigma,l)$>0. This concludes the proof.
Finite time consequences. Recalling that the finite time approximation $\sigma^{(n)}:=\lambda^{*n}*\delta_{\bf 0}$ converges to $\sigma$ exponentially fast, we may transfer the regularity properties of $\sigma$ to $\sigma^{(n)}$ provided we look at scales above an exponentially small threshold.
@@
    $\int_{\mathbb{R}^{d}}\lVert\boldsymbol{s}\rVert^{\gamma}\,\mathrm{d}\sigma^{(n)}(\boldsymbol{s})\ll 1,$
@@
    $\forall\varepsilon>e^{-n},\quad\sup_{\boldsymbol{s}\in\mathbb{R}^{d}}\sigma^{(n)}(\boldsymbol{s}+[-\varepsilon,\varepsilon]^{d})\ll\varepsilon^{\gamma}.$
Moreover, for $l\geq 1$, $c\lll_{l}1$, $P\in{\mathcal{P}}_{d,l}$ with $\|P\|=1$, $n\geq 1$, $\varepsilon>e^{-n}$, we have
  • (iii)
    $\sigma^{(n)}\{\boldsymbol{s}\in\mathbb{R}^{d}\,:\,|P(\boldsymbol{s})|\leq\varepsilon\}\ll_{l}\varepsilon^{c}.$
@@
In view of Lemma˜2.1, Items (i) and (ii) respectively follow from Lemma˜3.1 and Proposition˜3.2, see the proof of [7, Lemma 2.3] for more details. For item (iii), given $R>1$, $\varepsilon>0$, set $E_{R,\varepsilon}:=\{\boldsymbol{s}\in B_{R}\,:\,|P(\boldsymbol{s})|\leq\varepsilon\}$ where $B_{R}:=B^{\mathbb{R}^{d}}_{R}$. By Lemma˜2.1, we have for $1\ggg\varepsilon>e^{-n}$, for some $\gamma=\gamma(\lambda)\in(0,1)$,
@@
Observe $\|\nabla P\|_{B_{2R}}\ll R^{l}$. Therefore, taking $R=\varepsilon^{-\alpha}$ with $\alpha>0$, we have by ˜3.3
@@
where $c=c(\sigma,l)>0$. The result follows by taking $\alpha\lll_{l}1$, and applying Lemma˜3.5 (i) to allow restriction to $B_{\varepsilon^{-\alpha}}$. ∎
@@ 4. Effective recurrence of the $\mu$-walk
For $d=1$, a short self-contained proof is given in [7, Section 2.3]. For arbitrary $d$, we explain how to deduce Proposition˜4.1 from Prohaska-Sert-Shi [47], which is itself inspired by the works of Benoist-Quint [8], Eskin-Margulis [23], Eskin-Margulis-Mozes [24].
@@
(4.1) $k_{g}^{-1}a(\mathtt{r}_{g}^{-1})\in H_{\mu},\quad\quad\quad u(\mathtt{b}_{g})\in H_{\mu}.$
Write $S$ the set of vectors $\boldsymbol{s}\in\mathbb{R}^{d}$ such that $u(\boldsymbol{s})\in H_{\mu}$. Note $S$ is a Zariski-closed subgroup of $\mathbb{R}^{d}$, i.e. $S$ is a subspace. Using that $U\cap H_{\mu}$ is normalized by $H_{\mu}$, Equation (4.1), and the relation $a(\mathtt{r}_{g})k_{g}u(\boldsymbol{s})k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})=u(\mathtt{r}_{g}O_{g}\boldsymbol{s}+\mathtt{b}_{g})$, we get that $S$ is invariant under $\operatorname{supp}\lambda$. By irreducibility of $\lambda$, we deduce $S=\mathbb{R}^{d}$. This finishes the proof of Case ()(*).
@@
The measure $\mu^{\prime}:=\delta_{u(\boldsymbol{s}_{0})}*\mu*\delta_{u(-\boldsymbol{s}_{0})}$ on $P^{\prime}$ corresponds to some (strongly) irreducible randomized self-similar IFS $\lambda^{\prime}$. By the analysis of Case ()(*), we know that $H_{\mu^{\prime}}\supseteq U$. On the other hand, $H_{\mu^{\prime}}=u(\boldsymbol{s}_{0})H_{\mu}u(-\boldsymbol{s}_{0})$, whence $H_{\mu}\supseteq U$ as well. ∎
The previous lemma allows us to apply [47] to obtain some Margulis function, i.e. a proper positive function on $X$ which is uniformly contracted by the random walk and satisfies some growth control under the action of $G$.
@@
The combination of Lemma˜4.2 and [47, Corollary 3.8] guarantees that $\mu$ is $G$-expanding in the sense of [47, Definition 2.7]. This allows to apply [47, Theorem 6.1] to obtain the desired function $\beta$.
@@
For any proper function $\Upsilon:X\to[1,+\infty)$ satisfying properties (1) (2) of Lemma˜4.3, there exist $C,c>0$ such that for all $x\in X$,
@@
Using the assumptions on $\Upsilon$, the comparison between $\operatorname{inj}(x)$ and $\Upsilon(x)$ follows from the same argument as for [6, Lemmas 3.13, 3.14]. ∎
@@
Proof of Proposition˜4.1.
Let $\beta$ and $m,\theta,M$ as in Lemma˜4.3. We first replace $\beta$ by a suitable $\beta^{\prime}$ satisfying the contraction property with $m=1$. For that, given $f:X\rightarrow\mathbb{R}_{\geq 0}$, set $P_{\mu}f=\int_{G}f(g.)\,\mathrm{d}\mu(g)$. Let $\kappa>0$, consider $\beta^{\prime}=\beta+e^{\kappa}P_{\mu}\beta+\dots+e^{(m-1)\kappa}P_{\mu}^{m-1}\beta$. By Lemma˜4.3 (1), taking $\kappa:=\theta/m$, we have
(4.2) $P_{\mu}\beta^{\prime}\leq e^{-\kappa}\beta^{\prime}+e^{(m-1)\kappa}M.$
@@
where $M^{\prime}=e^{(m-1)\kappa}M/(1-e^{-\kappa})$. Using the Markov inequality, we deduce for every $\rho>0$,
@@
The proposition then follows from the comparison Lemma˜$4.$ ∎
As a direct corollary of Proposition˜4.1, we bound the Haar measure of cusp neighborhoods. This estimate will be useful in Section˜8.
@@
The Haar measure $m_{X}$ is an ergodic $\mu$-stationary measure. Hence for $m_{X}$-almost every $x\in X$, the sequence $(\mu^{*n}*\delta_{x})_{n\geq 0}$ converges to $m_{X}$ in Cesáro average for the weak-* topology. Then the first estimate follows immediately from Proposition˜4.1. Note the constants $C$, $c$ only depend on $\Lambda$ (and not $\mu$) because $\mu$ does not play a role in the statement.
By [6, Lemma 3.14], we have for $x\in X$,
(4.3) $\lvert\log\operatorname{inj}(x)\rvert-1\ll\operatorname{dist}(x,x_{0})\ll\lvert\log\operatorname{inj}(x)\rvert+1.$

### v2

@@ 4.2. Hölder regularity near subvarieties
We now bound $I_{3}$. We further decompose the integral according to the value of $\mathtt{r}_{\phi}$. By the choice of $n$ and the large deviation estimate (Crámer’s theorem) for $(\log\mathtt{r}_{\phi})_{\phi\sim\lambda^{*n}}$, there is $c_{3}=c_{3}(\lambda)>0$ such that
$\lambda^{*n}\bigl\{\,\phi\,:\,\mathtt{r}_{\phi}\notin[\varepsilon^{7/8},\varepsilon^{5/8}]\,\bigr\}\ll\varepsilon^{c_{3}}.$
Then
$I_{3}=\int_{\mathtt{b}_{\phi}\notin F_{\delta},\,\mathtt{r}_{\phi}\in[\varepsilon^{7/8},\,\varepsilon^{5/8}]}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi)\,+\,O(\varepsilon^{c_{3}}).$
In the next step, we remove from $E_{\varepsilon}$ anything outside the $\varepsilon^{1/2}$-ball centered at $\mathtt{b}_{\phi}=\phi(0)$. Indeed, for every $\phi$ satisfying $\mathtt{r}_{\phi}\leq\varepsilon^{5/8}$, we have
$\phi_{*}\sigma\bigl(\mathbb{R}^{d}\smallsetminus(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})\bigr)$ $=\sigma\bigl(\mathbb{R}^{d}\smallsetminus\phi^{-1}(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})\bigr)$
$=\sigma\bigl(\mathbb{R}^{d}\smallsetminus B_{\mathtt{r}_{\phi}^{-1}\varepsilon^{1/2}}\bigr)$
$\leq\sigma\bigl(\mathbb{R}^{d}\smallsetminus B_{\varepsilon^{-1/8}}\bigr).$
It follows then from Lemma˜3.1 that there is $c_{4}=c_{4}(\sigma)$ such that
$\int_{\mathtt{r}_{\phi}\leq\varepsilon^{5/8}}\phi_{\star}\sigma\bigl(\mathbb{R}^{d}\smallsetminus(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})\bigr)\,\mathrm{d}\lambda^{*n}(\phi)\leq\sigma\bigl(\mathbb{R}^{d}\smallsetminus B_{\varepsilon^{-1/8}}\bigr)\ll\varepsilon^{c_{4}}.$
Thus,
$I_{3}=\int_{\mathtt{b}_{\phi}\notin F_{\delta},\,\mathtt{r}_{\phi}\geq\varepsilon^{7/8}}\phi_{\star}\sigma\bigl(E_{\varepsilon}\cap(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})\bigr)\,\mathrm{d}\lambda^{*n}(\phi)+O(\varepsilon^{c_{3}}+\varepsilon^{c_{4}}).$
Consider $\phi$ as in the above integral and $\boldsymbol{s}\in E_{\varepsilon}\cap(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})$. By Taylor’s expansion and the fact that $\lVert\cdot\rVert_{C^{2}(B_{R+1})}\ll_{l}R^{l}\lVert\cdot\rVert$ on ${\mathcal{P}}_{d,l}$, the conditions $\|P\|=1$ and $\|\boldsymbol{s}-\mathtt{b}_{\phi}\|<\varepsilon^{1/2}$ imply
$\lvert P(\boldsymbol{s})-P(\mathtt{b}_{\phi})-\big\langle\nabla P(\mathtt{b}_{\phi}),\,\,\boldsymbol{s}-\mathtt{b}_{\phi}\big\rangle\rvert\ll_{l}\varepsilon R^{l}.$
As $\lvert P(\boldsymbol{s})\rvert\leq\varepsilon$, setting $v_{P,\phi}:=P(\mathtt{b}_{\phi})-\big\langle\nabla P(\mathtt{b}_{\phi}),\,\,\mathtt{b}_{\phi}\big\rangle$, we deduce
$\lvert\big\langle\nabla P(\mathtt{b}_{\phi}),\,\,\boldsymbol{s}\big\rangle-v_{P,\phi}\rvert\ll_{l}\varepsilon R^{l}$
But $\mathtt{b}_{\phi}\notin F_{\delta}$ means $\lVert\nabla P(\mathtt{b}_{\phi})\rVert>\delta$, whence $\boldsymbol{s}$ belongs to the $O_{l}(\varepsilon R^{l}\delta^{-1})$-neighborhood of an affine hyperplane. Applying $\phi^{-1}$ then the non-concentration near affine subspaces from Proposition˜4.2, and remembering $\mathtt{r}_{\phi}\geq\varepsilon^{7/8}$, we obtain
$\phi_{\star}\sigma\bigl(E_{\varepsilon}\cap(\mathtt{b}_{\phi}+B_{\varepsilon^{1/2}})\bigr)\ll_{l}(\mathtt{r}_{\phi}^{-1}\varepsilon R^{l}\delta^{-1})^{c_{5}}\leq(\varepsilon^{1/8}R^{l}\delta^{-1})^{c_{5}}$
where $c_{5}=c_{5}(\sigma)>0$.
We can conclude for $I_{3}$:
$I_{3}\ll_{l}(\varepsilon^{1/8}R^{l}\delta^{-1})^{c_{5}}+\varepsilon^{c_{3}}+\varepsilon^{c_{4}}.$ (4.11)
In the end, combining (4.8), (4.10), (4.11), and choosing $\delta=\varepsilon^{1/16}$, $R=\varepsilon^{-\alpha}$ with $\alpha\lll_{\sigma,l}1$, we have proven that for $c^{\prime}\lll_{\gamma,c_{1},\dots,c_{5}}1$,
$\sigma(E_{\varepsilon})\ll_{l}\varepsilon^{c^{\prime}}.$
@@
Let $k,l\geq 1$. Let $P:\mathbb{R}^{dk}\rightarrow\mathbb{R}$ be a polynomial map of degree at most $l$. Then for every $\varepsilon>0$,
$\sigma^{\otimes k}\bigl\{\,(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\,\leq\,\varepsilon\|P\|\,\bigr\}\leq C\varepsilon^{c}$
where $C,c>0$ depend only on $\sigma,k,l$.
@@
We consider $l$ fixed and argue by induction on $k$. The case $k=1$ is ˜4.4. Now given $k\geq 2$, we assume the result holds for k1k-1, with exponent $c=c(\sigma,k-1,l)$, and we prove it for $k$.
We may assume $\|P\|\geq 1$. We view $P$ as a polynomial in the indeterminates $(\boldsymbol{s}_{1},\dotsc,\boldsymbol{s}_{k-1})$ and with coefficients in the ring $\mathbb{R}[\boldsymbol{s}_{k}]$. Since $\lVert P\rVert\geq 1$, one of the coefficients $Q\in\mathbb{R}[\boldsymbol{s}_{k}]$ (of some monomial in $(\boldsymbol{s}_{1},\dotsc,\boldsymbol{s}_{k-1})$) satisfies $\lVert Q\rVert\geq 1$ (as an element of ${\mathcal{P}}_{d,l}$).
For any individual $\boldsymbol{s}_{k}\in\mathbb{R}^{d}$, the polynomial map
$\begin{array}[]{ccc}\mathbb{R}^{d(k-1)}&\to&\mathbb{R}\\ (\boldsymbol{s}_{1},\dotsc,\boldsymbol{s}_{k-1})&\mapsto&P(\boldsymbol{s}_{1},\dotsc,\boldsymbol{s}_{k-1},\boldsymbol{s}_{k})\end{array}$
seen as an element of ${\mathcal{P}}_{d(k-1),l}$ has norm at least $\lvert Q(\boldsymbol{s}_{k})\rvert$. Thus, by the induction hypothesis, whenever $\lvert Q(\boldsymbol{s}_{k})\rvert\geq\varepsilon^{1/2}$, we have
$\sigma^{\otimes(k-1)}\bigl\{\,(\boldsymbol{s}_{i})_{i=1}^{k-1}\in(\mathbb{R}^{d})^{k-1}\,:\,\lvert P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})\rvert\leq\varepsilon\,\bigr\}\ll_{k,l}\varepsilon^{c/2}$
We conclude by the Fubini-Lebesgue theorem and the case $k=1$ applied to $Q$:
$\sigma^{\otimes k}\bigl\{\,(\boldsymbol{s}_{i})_{i=1}^{k}$ $\in(\mathbb{R}^{d})^{k}\,:\,\lvert P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})\rvert\leq\varepsilon\,\bigr\}$
$=\int_{\mathbb{R}^{d}}\sigma^{\otimes(k-1)}\bigl\{\,(\boldsymbol{s}_{i})_{i=1}^{k-1}\in(\mathbb{R}^{d})^{k-1}\,:\,\lvert P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})\rvert\leq\varepsilon\,\bigr\}\,\mathrm{d}\sigma(\boldsymbol{s}_{k})$
$\ll_{k,l}\varepsilon^{c/2}+\sigma\bigl\{\,\boldsymbol{s}_{k}\in\mathbb{R}^{d}\,:\,\lvert Q(\boldsymbol{s}_{k})\rvert<\varepsilon^{1/2}\,\bigr\}$
$\ll_{k,l}\varepsilon^{c/2}+\varepsilon^{c(\sigma,1,l)/2}.$
This proves the case $k$ (with exponent $c(\sigma,k,l)=c(\sigma,k-1,l)/2$). ∎
4.3. Finite time consequences
Recalling that the finite time approximation $\sigma^{(n)}:=\lambda^{*n}*\delta_{0}$ converges to $\sigma$ exponentially fast, we may transfer the regularity properties of $\sigma$ to $\sigma^{(n)}$ provided we look at scales above an exponentially small threshold.
@@ 4.3. Finite time consequences
    $\forall\varepsilon>e^{-n},\qquad\sup_{\boldsymbol{s}\in\mathbb{R}^{d}}\sigma^{(n)}(\boldsymbol{s}+[-\varepsilon,\varepsilon]^{d})\ll\varepsilon^{\gamma}.$
Moreover, for $l\geq 1$, $c\lll_{l}1$, $P\in{\mathcal{P}}_{d,l}$ with $\lVert P\rVert=1$, $n\geq 1$, $\varepsilon>e^{-n}$, we have
@@
    $\sigma^{(n)}\bigl\{\,\boldsymbol{s}\in\mathbb{R}^{d}\,:\,\lvert P(\boldsymbol{s})\rvert\leq\varepsilon\,\bigr\}\ll_{l}\varepsilon^{c}.$
@@
In view of Lemma˜3.2, item (i) follows from Proposition˜4.2. For item (ii), given $R>1$, $\varepsilon>0$, set $E_{R,\varepsilon}:=\{\,\boldsymbol{s}\in B_{R}\,:\,\lvert P(\boldsymbol{s})\rvert\leq\varepsilon\,\}$ where $B_{R}:=B^{\mathbb{R}^{d}}_{R}$. By Lemma˜3.2, we have for $1\ggg\varepsilon>e^{-n}$, for some $\gamma=\gamma(\lambda)\in(0,1)$,
@@
Observe $\|\nabla P\|_{B_{2R}}\ll R^{l}$. Therefore, taking $R=\varepsilon^{-\alpha}$ with $\alpha>0$, we have by ˜4.4
@@
where $c=c(\sigma,l)>0$. The result follows by taking $\alpha\lll_{l}1$, and applying Lemma˜3.3 to allow restriction to $B_{\varepsilon^{-\alpha}}$. ∎
@@ 5. Effective recurrence of the $\mu$-walk
For $d=1$, a short self-contained proof is given in [3, Section 2.3]. For arbitrary $d$, we explain how to deduce Proposition˜5.1 from Prohaska-Sert-Shi [55], which is itself inspired by the works of Benoist-Quint [8], Eskin-Margulis [26], Eskin-Margulis-Mozes [25].
@@
$k_{g}^{-1}a(\mathtt{r}_{g}^{-1})\in H_{\mu},\quad\quad\quad u(\mathtt{b}_{g})\in H_{\mu}.$ (5.1)
Write $S$ the set of vectors $\boldsymbol{s}\in\mathbb{R}^{d}$ such that $u(\boldsymbol{s})\in H_{\mu}$. Note $S$ is a Zariski-closed subgroup of $\mathbb{R}^{d}$, i.e., $S$ is a subspace. Using that $U\cap H_{\mu}$ is normalized by $H_{\mu}$, Equation (5.1), and the relation $a(\mathtt{r}_{g})k_{g}u(\boldsymbol{s})k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})=u(\mathtt{r}_{g}O_{g}\boldsymbol{s}+\mathtt{b}_{g})$, we get that $S$ is invariant under $\operatorname{supp}\lambda$. By irreducibility of $\lambda$, we deduce $S=\mathbb{R}^{d}$. This finishes the proof of Case ()(*).
@@
The measure $\mu^{\prime}:=\delta_{u(\boldsymbol{s}_{0})}*\mu*\delta_{u(-\boldsymbol{s}_{0})}$ on $P^{\prime}$ corresponds to some randomized self-similar IFS $\lambda^{\prime}$. Namely, $\lambda^{\prime}$ is the pushfoward of $\lambda$ under conjugation by $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},v\mapsto v-s_{0}$; in particular $\lambda^{\prime}$ is irreducible with finite exponential moment. By the analysis of Case ()(*), we know that $H_{\mu^{\prime}}\supseteq U$. On the other hand, $H_{\mu^{\prime}}=u(\boldsymbol{s}_{0})H_{\mu}u(-\boldsymbol{s}_{0})$, whence $H_{\mu}\supseteq U$ as well. ∎
The previous lemma allows us to apply [55] to obtain some Margulis function, i.e., a proper positive function on $X$ which is uniformly contracted by the random walk and satisfies some growth control under the action of $G$.
@@
The combination of Lemma˜5.2 and [55, Corollary 3.8] guarantees that $\mu$ is $G$-expanding in the sense of [55, Definition 2.7]. This allows to apply [55, Theorem 6.1] to obtain the desired function $\beta$.
@@
For any proper function $\Upsilon:X\to[1,+\infty)$ satisfying properties (1) (2) of Lemma˜5.3, there exist $C,c>0$ such that for all $x\in X$,
@@
Using the assumptions on $\Upsilon$, the comparison between $\operatorname{inj}(x)$ and $\Upsilon(x)$ follows from the same argument as for [4, Lemmas 3.13, 3.14]. ∎
@@
Proof of Proposition˜5.1.
Let $\beta$ and $m,\theta,M$ as in Lemma˜5.3. We first replace $\beta$ by a suitable $\beta^{\prime}$ satisfying the contraction property with $m=1$. For that, let $\gamma,\kappa\in(0,1)$ be parameters to be specified below. Given $f:X\rightarrow\mathbb{R}_{\geq 0}$, set $P_{\mu}f=\int_{G}f(g\,\cdot)\,\mathrm{d}\mu(g)$. Consider $\beta^{\prime}=\beta^{\gamma}+e^{\kappa}P_{\mu}(\beta^{\gamma})+\dots+e^{(m-1)\kappa}P_{\mu}^{m-1}(\beta^{\gamma})$ and observe it is also proper.
Note that Lemma˜5.3 (2) implies $\beta^{\gamma}(gx)\leq\|\operatorname{Ad}(g)\|^{O(\gamma)}\beta^{\gamma}(x)$. Recalling $\mu$ has finite exponential moment and choosing $\gamma$ small enough accordingly, we obtain that $\beta^{\prime}$ takes finite values, and has controlled growth:
$\beta^{\prime}(gx)\leq\|\operatorname{Ad}(g)\|^{O(1)}\beta^{\prime}(x).$
Note also that $\beta^{\gamma}$ satisfies the contraction property of Lemma˜5.3 (1) with the parameters $\theta,M$ replaced by $(\gamma\theta,M^{\gamma})$. It follows that taking $\kappa:=\gamma\theta/m$, we have
$P_{\mu}\beta^{\prime}\leq e^{-\kappa}\beta^{\prime}+e^{(m-1)\kappa}M^{\gamma}.$ (5.2)
@@
where $M^{\prime}=e^{(m-1)\kappa}M^{\gamma}/(1-e^{-\kappa})$.
Using the Markov inequality, we deduce for every $\rho>0$,
@@
The proposition then follows from the comparison Lemma˜5.4 applied to $\beta^{\prime}$. ∎
As a direct corollary of Proposition˜5.1, we bound the Haar measure of cusp neighborhoods. This estimate will be useful in Section˜9.
@@
The Haar measure $m_{X}$ is an ergodic $\mu$-stationary measure. Hence for $m_{X}$-almost every $x\in X$, the sequence $(\mu^{*n}*\delta_{x})_{n\geq 0}$ converges to $m_{X}$ in Cesàro average with respect to the weak-* topology. Then the first estimate follows immediately from Proposition˜5.1. Note the constants $C$, $c$ only depend on $\Lambda$ (and not $\mu$) because $\mu$ does not play a role in the statement.
By [4, Lemma 3.14], we have for $x\in X$,
$\lvert\log\operatorname{inj}(x)\rvert-1\ll\operatorname{dist}(x,x_{0})\ll\lvert\log\operatorname{inj}(x)\rvert+1.$ (5.3)